# An improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$

### v2

@@ Abstract
We perform an improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$ ($L$ denotes the light ground state pseudoscalar, vector mesons and the corresponding $p$-wave scalar, axial-vector, and tensor ones) and predict their branching ratios (BRs) associated with relative ratios at leading order in the strong coupling $\alpha_{s}$. Our results ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})=(2.03^{+0.53}_{-0.41})\times 10^{-3}$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})=1.74^{+0.66}_{-0.50}$ are consistent with several available predictions in different approaches within uncertainties. Inputting the measured $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$ BRs with $p$ here being a proton, we derive the multibody $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ BRs through secondary decay chains via resonance $\eta_{c}$ under the narrow-width approximation, which might facilitate the (near) future tests of $B_{c}\to\eta_{c}$ decays. Under the $q\bar{q}$ assignment for light scalars, different to $B_{c}$ decaying into $J/\psi$ plus a scalar meson and other $B_{c}^{+}\to\eta_{c}L^{+}$ modes, surprisingly small $\Delta S=0$ BRs around ${\cal O}(10^{-7}-10^{-9})$ and highly large ratios near ${\cal O}(10^{2})$ between the $\Delta S=1$ and $\Delta S=0$ BRs are found in the $B_{c}$ decays to $\eta_{c}$ plus light scalars, with $S$ being strange number. Many large BRs and interesting ratios presented in this work could be tested by the Large Hadron Collider experiments, which would help us to examine the reliability of this improved perturbative QCD formalism for $B_{c}$-meson decays and further understand the QCD dynamics in the considered decay modes, as well as in the related hadrons.
@@ I Introduction
It is well known that, disparate from the $b\bar{b}$ and $c\bar{c}$ states, the $B_{c}$ meson is flavor-antisymmetric while unique since it is the only ground state containing two different heavy quarks $b$ and $c$ simultaneously Brambilla:2004wf ; Brambilla:2010cs . Peculiar to the more extensively studied $B_{u}$, $B_{d}$, and $B_{s}$ mesons, both constituents in a $B_{c}$ meson can decay separately, which offers a precious opportunity to expand our understanding of heavy $B$-meson physics through thoroughly exploring the more rich and complicated quantum chromodynamics (QCD) in the perturbative and nonperturbative regimes. Thus, the weak decays of the $B_{c}$ meson has prompted significant theoretical and experimental interest since its first discovery at Tevatron in 1998 CDF:1998axz ; CDF:1998ihx .
On the experimental side, the Large Hadron Collider (LHC) experiments starting running in 2009 have measured many $B_{c}$-meson decay channels of interest, such as $B_{c}^{+}\to J/\psi(\pi,2\pi,3\pi)$, even $B_{c}^{+}\to\chi_{cJ}\pi^{+}(J=0,1,2)$, etc. ParticleDataGroup:2024cfk ; HeavyFlavorAveragingGroupHFLAV:2024ctg . The $\eta_{c}$ meson, the lowest-lying $c\bar{c}$ pseudoscalar state, has attracted considerable theoretical and experimental attention since its discovery Partridge:1980vk . It decays primarily via $c\bar{c}$ annihilation into two gluons and is expected to have numerous hadronic decay modes into two- or three-body hadronic charged and/or neutral final states, which, unfortunately, seems not friendly for experimental studies at LHC experiments. Hence, though struggling against the large background and the small efficiency at LHC for these problems, the $B_{c}^{+}\to\eta_{c}\pi^{+}$ decay is not yet observed presently. However, along with the successful upgrade and resumed running of Large Hadron Collier-beauty (LHCb) detector since 2022, the ensuingly exciting results are forthcoming Gao:2024lhc . The continuously collected data with this accomplished even the future prospective upgrades of LHCb detector certainly will offer us a precious opportunity to promote the $B_{c}$-meson physics into a precision era. It is therefore expected that a huge amount of data with superior quality can facilitate a promising measurement of $B_{c}^{+}\to\eta_{c}\pi^{+}$, for example, via $\eta_{c}\to p\bar{p}$ decay chain. Here, $p$ denotes a proton.
@@
Table 1: Various predictions for ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$, ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}K^{+})/{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ in the literature.
Observables Chang:1992pt Anisimov:1998uk Colangelo:1999zn AbdElHady:1999xh Verma:2001hb Ebert:2003cn Ivanov:2006ni Hernandez:2006gt Choi:2009ym  [^1] The results are calculated on the basis of Coulomb plus linear confining and harmonic oscillator (in the parentheses) potentials. Naimuddin:2012dy Qiao:2012hp  [^2] These results are calculated with the non-relativistic QCD approach at leading order in the strong coupling $\alpha_{s}$. And the corresponding next-to-leading order BR in $\alpha_{s}^{2}$ is $(5.19^{+0.70}_{-1.07})\times 10^{-3}$. Rui:2014tpa Issadykov:2018myx Nayak:2022qaq Wu:2024gcq Deng:2025znr  [^3] These results are calculated with the QCD factorization approach in $\alpha_{s}$. And the corresponding next-to-next-to-leading order BR in $\alpha_{s}^{2}$ is $(0.81^{+0.88}_{-0.56})\times 10^{-3}$.
$10^{3}\cdot{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ $1.80$ $1.30$ $0.26$ $1.40$ $9.30$ $0.85$ $1.90$ $0.94$ $0.91(1.16)$ $0.34$ $2.95$ $2.98^{+1.24}_{-1.05}$ $1.89^{+0.37}_{-0.37}$ $0.397$ $1.40^{+0.40}_{-0.40}$ $0.79^{+0.86}_{-0.55}$
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $1.00$ $1.78$ $0.19$ $1.27$ $2.07$ $1.39$ $1.12$ $1.24$ $\dots$ $1.00$ $1.33$ $1.28^{+0.69}_{-0.56}$ $1.87^{+0.52}_{-0.52}$ $1.02$ $0.56^{+0.21}_{-0.21}$ $0.95^{+1.04}_{-0.68}$
$10^{2}\cdot\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}$ $7.78$ $10.00$ $8.00$ $7.86$ $5.05$ $8.24$ $7.89$ $7.98$ $8.13(8.10)$ $8.82$ $7.12$ $8.05^{+0.68}_{-0.99}$ $7.94^{+0.02}_{-0.05}$ $7.81$ $7.86^{+0.14}_{-0.08}$ $7.51^{+0.02}_{-0.00}$
Table 2: Same as Table 1 but for $B_{c}^{+}\to\eta_{c}(\rho,K^{*})^{+}$.
Observables Chang:1992pt Anisimov:1998uk Colangelo:1999zn AbdElHady:1999xh Verma:2001hb Ebert:2003cn Ivanov:2006ni Hernandez:2006gt Choi:2009ym Naimuddin:2012dy Qiao:2012hp Rui:2014tpa Issadykov:2018myx Nayak:2022qaq Wu:2024gcq Deng:2025znr
$10^{3}\cdot{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})$ $4.90$ $3.00$ $0.67$ $3.30$ $3.70$ $2.10$ $4.50$ $2.40$ $2.57(3.24)$ $1.06$ $7.89$ $9.83^{+3.33}_{-2.59}$ $5.18^{+1.04}_{-1.04}$ $1.24$ $3.80^{+0.10}_{-0.10}$ $2.15^{+2.23}_{-1.46}$
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $2.72$ $4.11$ $0.52$ $3.0$ $82.2$ $3.44$ $2.65$ $3.16$ $\dots$ $3.12$ $3.55$ $4.22^{+2.05}_{-1.57}$ $5.13^{+1.45}_{-1.45}$ $3.18$ $1.52^{+0.37}_{-0.37}$ $2.60^{+2.71}_{-1.80}$
$10^{2}\cdot\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{*+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}$ $5.10$ $7.00$ $5.97$ $5.45$ $3.70$ $5.24$ $5.56$ $5.42$ $5.06(5.25)$ $5.66$ $5.20$ $5.80^{+0.52}_{-0.45}$ $5.60^{+0.03}_{-0.04}$ $5.24$ $5.53^{+1.14}_{-1.48}$ $5.12^{+0.15}_{-0.06}$
@@
To the best of our knowledge, the nature of light scalars, especially those under or near 1 GeV, remains a long-standing puzzle in hadron physics. Recently, the CMS Collaboration found strong evidence of $f_{0}(980)$ being a normal quark-antiquark state CMS:2023rev , even though the ALICE Collaboration supported the $K_{0}^{*}(700)$ being a four-quark state ALICE:2023eyl . Undoubtedly, more endeavors need to be devoted to this field. The useful clues about the nature of light scalars could be collected indirectly through probing their productions in the heavy hadron decays, just like the $B\to f_{0}(980)K$ decays observed in the $B$-factory experiments Abe:2002av ; Aubert:2003mi . So far, there are two different scenarios to describe these scalar mesons $a_{0}(980,1450)$ and $K_{0}^{*}(700,1430)$ under the $q\bar{q}$ assignment Cheng:2005nb . In scenario 1 ($S1$), the scalar mesons $a_{0}(980)$ and $K_{0}^{*}(700)$ are treated as the lowest-lying states, and those $a_{0}(1450)$ and $K_{0}^{*}(1430)$ are the first excited states correspondingly. And, in scenario 2 ($S2$), the scalar mesons $a_{0}(1450)$ and $K_{0}^{*}(1430)$ are viewed as the ground states, while those $a_{0}(980)$ and $K_{0}^{*}(700)$ might be the four-quark states. However, as stressed in Cheng:2005nb , it is difficult in practice to make quantitative predictions based on the four-quark or tetraquark picture for light scalars because the calculations of decay constant and form factors of light scalars are beyond the conventional quark model and the involved nonfactorizable contributions cannot be calculated in the available QCD-based factorization framework. Moreover, the productions of light scalars from the vacuum in the $B_{c}^{+}\to\eta_{c}S^{+}$ decays are expected to be highly suppressed originating from the nearly-zero vector decay constants $f_{S}$ ( Actually, the vector decay constants $f_{S}=0$ in the $SU(3)$ limit.) Cheng:2005nb . It means that investigations on these decays must go beyond naive factorization. Therefore, the predictions in this work are made on the basis of two-quark model for light scalars within the iPQCD framework.
@@
The $p$-wave light axial-vectors have been investigated at both experimental and theoretical aspects. However, our understanding about their nature is still far from complete Du:2022nno . In the spectroscopy study ParticleDataGroup:2024cfk , $a_{1}(1260)$ and $b_{1}(1235)$ are the $1^{3}\!P_{1}$ and $1^{1}\!P_{1}$ axial-vector states, respectively, carrying quantum numbers $J^{PC}=1^{++}$ and $J^{PC}=1^{+-}$ correspondingly. While it is very interesting that the strange $K_{1}(1270)$ and $K_{1}(1400)$ mesons are generally regarded as the mixtures of $1^{3}\!P_{1}$ state $K_{1A}$ and $1^{1}\!P_{1}$ state $K_{1B}$ due to the $SU(3)$ flavor broken symmetry Cheng:2011pb ,
@@
with mixing angle $\theta_{K}$. The value of $\theta_{K}$ can be related to the masses of the $K_{1}(1270)$ and $K_{1}(1400)$, to the strong decays of the $K_{1}(1270)$ and $K_{1}(1400)$, and to rates of weak decays to final states involving the $K_{1}(1270)$ and $K_{1}(1400)$ Suzuki:1993yc ; Blundell:1995au . Thus, the decays such as $B_{c}^{+}\to(c\bar{c})K_{1}(1270,1400)^{+}$ would be of great interest for exploring the information of $\theta_{K}$. However, presently, there is no consensus on the value of the mixing angle $\theta_{K}$, and the results from various approaches are still quite controversial, e.g., see a short overview in Liu:2024lph ( and references therein). We therefore take both referenced values, i.e., $\theta_{K_{1}}\approx 33^{\circ}$ and $58^{\circ}$ Cheng:2011pb ; Shi:2023kiy into account in the related numerical calculations of this work.
@@
By incorporating the finite charm quark mass effects into Sudakov resummation of the large logarithmic corrections to wave functions through $k_{T}$ resummation at the next-to-leading-logarithm accuracy Liu:2020upy , besides including them in the hard kernel, the iPQCD formalism is now self-consistent for systematically studying the $B_{c}$-meson decays and $B$-meson decaying into charmonia. So far, facilitated by the newly proposed transverse-momentum-dependent $B_{c}$-meson wave function Liu:2018kuo , we have studied the $B_{c}^{+}\to J/\psi M^{+}$ Liu:2023kxr and $B_{c}^{+}\to\chi_{cJ}(P,V)^{+}$ Liu:2025ipe decays, in which, the BRs predicted in the iPQCD formalism are generally consistent with several available predictions in other approaches. Furthermore, the resultant relative ratios such as ${\rm BR}(B_{c}^{+}\to J/\psi a_{1}(1260)^{+}(\to\pi^{+}\pi^{-}\pi^{+}))/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, ${\rm BR}(B_{c}^{+}\to J/\psi\rho^{+}(\to\pi^{+}\pi^{0}))/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, ${\rm BR}(B_{c}^{+}\to\chi_{c2}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ etc. agree well with the current data reported by the LHC experiments within theoretical uncertainties. In light of these successful output, we shall analyze the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism at leading order in the strong coupling $\alpha_{s}$ Liu:2018kuo ; Liu:2020upy . Inputting the BRs ParticleDataGroup:2024cfk of strong decays $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$, we additionally present the BRs of multibody modes arising from $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ via resonance $\eta_{c}$ under the narrow-width approximation. The (near) future tests of our iPQCD predictions at experimental facilities are expected to help us to further understand the QCD dynamics involved in these $B_{c}$-meson decays, even explore the inner structure of $\eta_{c}$ and related light hadrons.
@@
The rest of this paper is organized as follows. In Sect. II, the formalism and the perturbative calculations in association with factorization formulas of $B_{c}^{+}\to\eta_{c}L^{+}$ are presented. The numerical results and phenomenological analyses are given in Sect. III. Sect. IV summarizes our main conclusions.

### v3

# Improved perturbative QCD study of the decay $B_{c}^{+}\to\eta_{c}L^{+}$

@@ Abstract
We perform an improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$, where $L$ denotes the light ground-state pseudoscalar, vector mesons and the corresponding $p$-wave scalar, axial-vector, and tensor ones, and predict their branching ratios (BRs) associated with relative ratios at leading order in the strong coupling $\alpha_{s}$. Our results ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})=(2.03^{+0.53}_{-0.41})\times 10^{-3}$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})=1.74^{+0.66}_{-0.50}$ are consistent with several available predictions in different approaches within uncertainties. Inputting the measured $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$ BRs with $p$ here being a proton, we derive the multibody $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ BRs through secondary decay chains via resonance $\eta_{c}$ under the narrow-width approximation, which might facilitate the (near-) future tests of $B_{c}\to\eta_{c}$ decays. Under the $q\bar{q}$ assignment for light scalars, different to $B_{c}$ decaying into $J/\psi$ plus a scalar meson and other $B_{c}^{+}\to\eta_{c}L^{+}$ modes, surprisingly small $\Delta S=0$ BRs around ${\cal O}(10^{-7}-10^{-9})$ and highly large ratios near ${\cal O}(10^{2})$ between the $\Delta S=1$ and $\Delta S=0$ BRs are found in the $B_{c}$ decays to $\eta_{c}$ plus light scalars, with $S$ being strange number. Many large BRs and interesting ratios presented in this work could be tested by the Large Hadron Collider experiments, which would help us to examine the reliability of this improved perturbative QCD formalism for $B_{c}$-meson decays and further understand the QCD dynamics in the considered decay modes as well as in the related hadrons.
@@ I Introduction
It is well known that, disparate from the $b\bar{b}$ and $c\bar{c}$ states, the $B_{c}$ meson is flavor antisymmetric while unique since it is the only ground state containing two different heavy quarks $b$ and $c$ simultaneously Brambilla:2004wf ; Brambilla:2010cs . Peculiar to the more extensively studied $B_{u}$, $B_{d}$, and $B_{s}$ mesons, both constituents in a $B_{c}$ meson can decay separately, which offers a precious opportunity to expand our understanding of heavy $B$-meson physics through thoroughly exploring the more rich and complicated QCD in the perturbative and nonperturbative regimes. Thus, the weak decays of the $B_{c}$ meson has prompted significant theoretical and experimental interest since its first discovery at the Tevatron in 1998 CDF:1998axz ; CDF:1998ihx .
On the experimental side, the LHC experiments that started running in 2009 have observed many $B_{c}$-meson decay channels of interest, such as $B_{c}^{+}\to J/\psi(\pi,2\pi,3\pi)^{+}$, even $B_{c}^{+}\to\chi_{cJ}\pi^{+}(J=0,1,2)$, etc. ParticleDataGroup:2024cfk ; HeavyFlavorAveragingGroupHFLAV:2024ctg . The $\eta_{c}$ meson, the lowest-lying $c\bar{c}$ pseudoscalar state, has attracted considerable theoretical and experimental attention since its discovery Partridge:1980vk . It decays primarily via $c\bar{c}$ annihilation into two gluons and is expected to have numerous hadronic decay modes into two- or three-body hadronic charged and/or neutral final states, which, unfortunately, seems not friendly for experimental studies at LHC experiments. Hence, though struggling against the large background and the small efficiency at LHC for these problems, the $B_{c}^{+}\to\eta_{c}\pi^{+}$ decay is not yet observed presently. However, along with the successful upgrade and resumed running of Large Hadron Collier-beauty (LHCb) detector since 2022, the ensuingly exciting results are forthcoming Gao:2024lhc . The continuously collected data from the upgraded LHCb detector will offer a precious opportunity to advance $B_{c}$-meson physics into a precision era. It is therefore expected that a huge amount of data with superior quality can facilitate a promising measurement of $B_{c}^{+}\to\eta_{c}\pi^{+}$, for example, via $\eta_{c}\to p\bar{p}$ decay chain. Here, $p$ denotes a proton.
@@
To the best of our knowledge, the nature of light scalars, especially those under or near 1 GeV, remains a long-standing puzzle in hadron physics. Recently, the CMS Collaboration found strong evidence of $f_{0}(980)$ being a normal quark-antiquark state CMS:2023rev , even though the ALICE Collaboration supported the $K_{0}^{*}(700)$ being a four-quark state ALICE:2023eyl . Undoubtedly, more endeavors need to be devoted to this field. The useful clues about the nature of light scalars could be collected indirectly through probing their productions in the heavy hadron decays, just like the $B\to f_{0}(980)K$ decays observed in the $B$-factory experiments Abe:2002av ; Aubert:2003mi . So far, there are two different scenarios to describe these scalar mesons $a_{0}(980,1450)$ and $K_{0}^{*}(700,1430)$ under the $q\bar{q}$ assignment Cheng:2005nb . In scenario 1 ($S1$), the scalar mesons $a_{0}(980)$ and $K_{0}^{*}(700)$ are treated as the lowest-lying states, and those $a_{0}(1450)$ and $K_{0}^{*}(1430)$ are the first excited states correspondingly. And, in scenario 2 ($S2$), the scalar mesons $a_{0}(1450)$ and $K_{0}^{*}(1430)$ are viewed as the ground states, while those $a_{0}(980)$ and $K_{0}^{*}(700)$ might be the four-quark states. However, as stressed in Ref. Cheng:2005nb , it is difficult in practice to make quantitative predictions based on the four-quark or tetraquark picture for light scalars because the calculations of decay constant and form factors of light scalars are beyond the conventional quark model and the involved nonfactorizable contributions cannot be calculated in the available QCD-based factorization framework. Moreover, the productions of light scalars from the vacuum in the $B_{c}^{+}\to\eta_{c}S^{+}$ decays are expected to be highly suppressed originating from the nearly zero vector decay constants $f_{S}$ [ actually, the vector decay constants $f_{S}=0$ in the $SU(3)$ limit.] Cheng:2005nb . It means that investigations on these decays must go beyond naive factorization. Therefore, the predictions in this work are made on the basis of two-quark model for light scalars within the iPQCD framework.
@@
The $p$-wave light axial vectors have been investigated at both experimental and theoretical aspects. However, our understanding about their nature is still far from complete Du:2022nno . In the spectroscopy study ParticleDataGroup:2024cfk , $a_{1}(1260)$ and $b_{1}(1235)$ are the $1^{3}\!P_{1}$ and $1^{1}\!P_{1}$ axial-vector states, respectively, carrying quantum numbers $J^{PC}=1^{++}$ and $J^{PC}=1^{+-}$ correspondingly. Although it is very interesting that the strange $K_{1}(1270)$ and $K_{1}(1400)$ mesons are generally regarded as the mixtures of $1^{3}\!P_{1}$ state $K_{1A}$ and $1^{1}\!P_{1}$ state $K_{1B}$ due to the $SU(3)$ flavor broken symmetry Cheng:2011pb ,
@@
with mixing angle $\theta_{K}$. The value of $\theta_{K}$ can be related to the masses of the $K_{1}(1270)$ and $K_{1}(1400)$, to the strong decays of the $K_{1}(1270)$ and $K_{1}(1400)$, and to rates of weak decays to final states involving the $K_{1}(1270)$ and $K_{1}(1400)$ Suzuki:1993yc ; Blundell:1995au . Thus, the decays such as $B_{c}^{+}\to(c\bar{c})K_{1}(1270,1400)^{+}$ would be of great interest for exploring the information of $\theta_{K}$. However, presently, there is no consensus on the value of the mixing angle $\theta_{K}$, and the results from various approaches are still quite controversial; e.g., see a short overview in Ref. Liu:2024lph ( and references therein). We therefore take both referenced values, i.e., $\theta_{K_{1}}\approx 33^{\circ}$ and $58^{\circ}$ Cheng:2011pb ; Shi:2023kiy into account in the related numerical calculations of this work.
@@
By incorporating the finite charm quark mass effects into Sudakov resummation of the large logarithmic corrections to wave functions through $k_{T}$ resummation at the next-to-leading-logarithm accuracy Liu:2020upy , besides including them in the hard kernel, the iPQCD formalism is now self-consistent for systematically studying the $B_{c}$-meson decays and $B$-meson decaying into charmonia. So far, facilitated by the newly proposed transverse-momentum-dependent $B_{c}$-meson wave function Liu:2018kuo , we have studied the $B_{c}^{+}\to J/\psi M^{+}$ Liu:2023kxr and $B_{c}^{+}\to\chi_{cJ}(P,V)^{+}$ Liu:2025ipe decays, in which, the BRs predicted in the iPQCD formalism are generally consistent with several available predictions in other approaches. Furthermore, the resultant relative ratios such as ${\rm BR}(B_{c}^{+}\to J/\psi a_{1}(1260)^{+}(\to\pi^{+}\pi^{-}\pi^{+}))/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, ${\rm BR}(B_{c}^{+}\to J/\psi\rho^{+}(\to\pi^{+}\pi^{0}))/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, ${\rm BR}(B_{c}^{+}\to\chi_{c2}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, etc. agree well with the current data reported by the LHC experiments within theoretical uncertainties. In light of these successful outputs, we shall analyze the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism at leading order in the strong coupling $\alpha_{s}$ Liu:2018kuo ; Liu:2020upy . Inputting the BRs ParticleDataGroup:2024cfk of strong decays $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$, we additionally present the BRs of multibody modes arising from $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ via resonance $\eta_{c}$ under the narrow-width approximation. The (near-)future tests of our iPQCD predictions at experimental facilities are expected to help us to further understand the QCD dynamics involved in these $B_{c}$-meson decays, and even explore the inner structure of $\eta_{c}$ and related light hadrons.
@@
The rest of this paper is organized as follows. In Sec. II, the formalism and the perturbative calculations in association with factorization formulas of $B_{c}^{+}\to\eta_{c}L^{+}$ are presented. The numerical results and phenomenological analyses are given in Sec. III. Section IV summarizes our main conclusions.